OpenStudy (anonymous):

evaluate the integral (√(1/x)-2) dx from 1 to 4

OpenStudy (raden):

is the -2 in the square root or not ?

OpenStudy (anonymous):

it is not

OpenStudy (raden):

sorry, i have to go now. let me tag my best friends :) @ganeshie8 , @whpalmer4

OpenStudy (whpalmer4):

So is that\[\int_1^4 (\sqrt{\frac{1}{x}} -2) dx\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Well, we can split this into two integrals: \[\int_1^4 \sqrt{\frac{1}{x}}\,dx - \int_1^4 2\,dx \]right? Can you tell me the result for the right-hand integral?

OpenStudy (whpalmer4):

The left hand integral you can rewrite if you recall that \[\large \sqrt{\frac{a}{b}} = \frac{\sqrt{a}}{\sqrt{b}}\]and \[\large \sqrt{x} = x^{\frac{1}{2}}\]

OpenStudy (whpalmer4):

That gives you an expression to which you can apply the power rule for integration.